# MagicDriveDiT: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v1

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is essential for applications like autonomous driving. However, existing methods are limited by scalability and how control conditions are integrated, failing to meet the needs for high-resolution and long videos for autonomous driving applications. In this paper, we introduce MagicDriveDiT, a novel approach based on the DiT architecture, and tackle these challenges. Our method enhances scalability through flow matching and employs a progressive training strategy to manage complex scenarios. By incorporating spatial-temporal conditional encoding, MagicDriveDiT achieves precise control over spatial-temporal latents. Comprehensive experiments show its superior performance in generating realistic street scene videos with higher resolution and more frames. MagicDriveDiT significantly improves video generation quality and spatial-temporal controls, expanding its potential applications across various tasks in autonomous driving.
@@ 1 Introduction
With the rapid development of diffusion models, video synthesis [3, 14, 1, 2] has seen significant advancements. Controllable video synthesis [20, 16] greatly facilitates the use of synthetic data in numerous downstream applications [37]. Especially in the field of autonomous driving, technologies like MagicDrive [11] and DriveDreamer [35] exemplify how controllable multi-view video synthesis, capable of generating high-quality street-view videos, effectively aid downstream tasks, and enhance the performance and reliability of autonomous systems.
High-resolution and long video synthesis is a focal point of research in the content generation domain [27, 5], and is highly demanded by autonomous driving technology. Specifically, autonomous driving perception models require higher-resolution inputs to discern details and distant objects within the data [22]. The application of video synthesis in 3D reconstruction also necessitates high resolution to capture intricate details [10, 45]. On the other hand, long videos provide more content for more comprehensively evaluating autonomous driving algorithms [42]. From a technical perspective, high resolution is a straightforward method to enhance generation quality [5], while long videos encompass more dynamic content [40], aiding models in effectively learning the laws of the world [3].
@@
However, synthesizing high-resolution long videos presents two major challenges. Firstly, the increase in data volume require the generative model’s high scalability to support high-quality synthesis effectively. Secondly, geometry control in video generation poses significant challenges. Previous controllable techniques are primarily designed for image generation [11, 35, 39, 38, 27, 18], focusing on spatial latent spaces. Achieving high-resolution long video synthesis necessitates a 3D compressed VAE that integrates spatial and temporal information [3, 47, 43, 19]. Utilizing spatial-temporal latents renders previous control methods inadequate, as demonstrated in Figure 2. This necessitates novel approaches to effectively manage elements in synthetic videos, ensuring both spatial and temporal coherence with the given conditions.
@@
In this paper, we introduce MagicDriveDiT, a novel approach based on the DiT architecture, designed to address the above challenges for high-resolution and long street-view video synthesis with precise control. We adopt the flow matching [9] formulation for diffusion models, implementing it with a DiT-based architecture [30, 47], to enhance scalability. This improves the model’s capacity to efficiently handle complex data across diverse scenarios. To enhance controllability, we use specially designed spatial-temporal conditional encoding for spatial-temporal latents from CogVAE [43], enabling precise management of representations. Besides, we adopt a progressive bootstrapping strategy for model training, transitioning from short to long videos, allowing the model to capture intricate details and generalize to complex scenarios. Additionally, we utilize videos of various resolutions and durations to strengthen the model’s generalization ability, enabling the model to synthesize longer videos than those for training.
Our MagicDriveDiT excels in generating highly realistic videos that align with road maps, 3D bounding boxes, and different camera perspectives, achieving higher resolution and more frames than previous works [11, 12, 27, 18]. Comprehensive experiments and comparisons demonstrate the effectiveness of our training and control methods, significantly improving controllable street view video synthesis. MagicDriveDiT’s flexibility in handling various resolutions, frames, and control signals enables the creation of novel street views suitable for simulations, expanding its potential applications across diverse domains.
@@
  • We design an efficient framework, MagicDriveDiT, leveraging progressive bootstrapping to achieve high-quality high-resolution long video generation.
  • We develop novel spatiotemporal control for object positions, road semantics, and camera trajectories while maintaining multi-frame, multi-view consistency.
@@
Figure 3: Architecture Overview of MagicDriveDiT. To incorporate different conditions for video generation, MagicDriveDiT adopts a two-branch architecture as in [44] with basic STDiT3 blocks from [6, 47]. We also propose the MVDiT block for multi-view consistency and Spatial-Temporal (Box/Traj.) Encoder to inject condition into the Spatial-Temporal (SP) Latent.
@@ 2 Related Work
Video Generation in Autonomous Driving. Video generation is crucial for autonomous driving, with application in training perception models [11], testing [42], and scene reconstruction [10, 45]. It demands extensive fields of view and dynamic object motion handling, requiring generation model controllability [11] and high-resolution video production [22] with more frames [10] and multiple camera perspectives [45]. Diffusion models have improved controllable multi-view video generation, yet existing models [38, 39, 46, 27, 18] lack adequate resolution and frame count for data engine applications [11, 10, 45] and policy testing [17]. As compared in Table 1, our work, MagicDriveDiT, achieves unprecedented resolution and frame count in video generation for autonomous driving.
@@
MagicDriveDiT 848$\times$1600$\times$6 129
424$\times$800$\times$6 241
Table 1: Comparison of Resolution and Frame Count. We only consider a single inference, since rollout notably degrades quality. Only support text & image(s) conditions. See note in Appendix.
Diffusion Models and DiT Architectures. Diffusion models [13, 32, 48] generate data by learning the denoising steps from Gaussian noises to samples, widely used in image [8, 31, 37] and video generation [14]. From the modeling’s perspective, Flow Matching [23, 9] simplifies the modeling of diffusion models, enhancing the efficiency of training and inference. From the implementation’s perspective, the architecture of diffusion models shifts from UNet [13] to DiT, due to better scalability [30, 6], especially for high-resolution tasks [5]. Our MagicDriveDiT also leverages Flow Matching and DiT for scaling to high-resolution and long video generation.
Conditional Generation. Conditional generation is crucial for various applications utilizing generative models. Cross-attention layers from LDM [31] and ControlNet’s [44] additive encoding for grid-shaped control signals are leading methods in controllable diffusion-based generation. In street view generation, MagicDrive [11] and MagicDrive3D [10] integrate 3D bounding boxes, BEV maps, ego trajectories, and camera poses for multi-view street scene synthesis. However, these methods are limited to spatial encoding and are not directly applicable to spatial-temporal VAE latents [43], as shown in Figure 2. MagicDriveDiT presents a novel control paradigm for spatiotemporally compressed VAE latents, enabling the controllable generation of high-resolution long videos.
@@ 3 Preliminary
Problem Formulation. This paper addresses controllable high-resolution, long video generation for street views. Given a sequence of frame descriptions $\{\mathbf{S}_{t}\},t\in\{0,\dots,T\}$, the goal is to generate corresponding street-view videos from latent variables $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$, i.e., $\{\mathbf{I}_{c,t}\}=\mathcal{G}(\{\mathbf{S}_{t}\},\mathbf{z})$, where $c\in\{0,\dots,C\}$ denotes the $C$ camera views. Here, high resolution indicates that $\mathbf{I}$ has a high resolution, and long video implies that $T$ is large.
To describe a street-view video, we adopt the conditions outlined in [10, 11]. Specifically, frame descriptions $\mathbf{S}_{t}=\{\mathbf{C},\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L},\mathbf{Tr}_{t}^{0}\}$ include the camera poses $\{\mathbf{C}_{c}\}=[\mathbf{R}_{c},\mathbf{t}_{c}]$ [^1] Typically, cameras are fixed throughout the video., a road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ representing a $w\times h$ meter road area in BEV with $c$ semantic classes, 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ where each object is described by a box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$, text $\mathbf{L}$ adding information for the whole video (e.g., weather and time of day), and the ego vehicle trajectory $\mathbf{Tr}_{t}^{0}$ which describes the transformation from the LiDAR coordinate of each frame to the first frame, i.e., $\mathbf{Tr}^{0}_{t}=[\mathbf{R}_{t}^{0},\mathbf{t}_{t}^{0}]$. All geometric information, except for $\mathbf{Tr}_{t}^{0}$, is parameterized according to the local LiDAR coordinate of the ego car.
@@ 4.1 Overview of MagicDriveDiT
For model architecture, depicted in Figure 3, MagicDriveDiT introduces a novel DiT-based diffusion model for controllable street-view generation, utilizing STDiT-3 blocks as per Zheng et al. [47]. These blocks feature separate spatial and temporal modules to enhance information processing.
The architecture design incorporates two significant modifications. First, to facilitate multi-view generation, the Multi-View DiT (MVDiT) block integrates a cross-view attention layer [11], as shown on the left of Figure 3. Second, given the need for handling multiple control elements, MagicDriveDiT employs cross-attention [31] for text, boxes, camera views, and trajectories, alongside an addictive branch [44] for maps to inject control signals.
@@
Furthermore, MagicDriveDiT employs a progressive bootstrap training strategy (Section 4.3), transitioning from images to low-resolution and short videos, and ultimately to high-resolution and long videos, to enhance diffusion model convergence. The final training stage incorporates variable length and resolution videos, allowing the model to generate images and videos at diverse resolutions and extrapolate to longer frame counts beyond the training settings. Section 4.4 provides further details.
4.2 Design for High-Resolution Long Video
DiT and 3D VAE. Training diffusion models for high-resolution and long video generation is computationally intensive and requires substantial GPU memory. DiT and 3D VAE are pivotal in scaling diffusion models in these aspects.
As noted by Peebles and Xie [30], there is a strong negative correlation between model Gflops and FID, making DiT preferable to architectures like UNet [11] for computational efficiency. For GPU memory, 3D VAE offers temporal downsampling. A typical 2D VAE [31] compresses images by a factor of 8 in both height and width, achieving a 64×64$\times$64 × compression ratio. In contrast, 3D VAE [43] compresses temporal information by 4×4$\times$4 ×, resulting in a 256×256$\times$256 × compression ratio [^2] We ignore possible discrepancy in the dimension of latents through our discussion., significantly reducing the sequence length of pachified latent and memory consumption, particularly beneficial for transformers [7]. Additionally, the trend of architectural unification [34] allows DiT to utilize advanced parallelization methods, e.g., sequence parallelism [47], to overcome single GPU memory constraints, facilitating higher resolution and longer video generation.
A major challenge with 3D VAE adoption is geometric control. As illustrated in Figure 2, geometric control manages per-frame content spatially. With 2D VAE, $T$ frames are encoded into $T$ latents. Using $T$ geometric descriptors $\{\mathbf{S}_{t}\},t\in\{1,...,T\}$, video geometric control degrades to spatial control on images [28], as control signals and latents are temporally aligned. However, 3D VAE yields $T/f$ latents (where $f$ is the temporal compression ratio), misaligning control signals with latents and rendering previous control techniques [11, 27, 39] ineffective.
@@ 4.2 Design for High-Resolution Long Video
Spatial-Temporal Conditioning Techniques. For geometric control in MagicDriveDiT, we introduce spatial-temporal encoding to align control signals with spatial-temporal latents. This involves re-aligning the maps ($\mathbf{M}_{t}$), boxes ($\mathbf{B}_{t}$), and trajectory ($\mathbf{Tr}_{t}^{0}$) within the scene descriptor ($\mathbf{S}_{t}$). Maps, represented as grid data, are straightforward to manage. By extending ControlNet’s [44] design, we utilize temporal downsampling modules with new trainable parameters in 3D VAE to align features between control and base blocks, as depicted in Figure 4(a).
@@
For 3D boxes, padding is applied to invisible boxes to maintain consistent box sequence length across views and frames, illustrated in Figure 3 right. We employ a downsampling module with a temporal transformer and RoPE [33] to capture temporal correlation, creating spatial-temporal embeddings aligned with video latents, as in Figure 4(b). The spatial-temporal encoder for boxes can also be adapted for ego trajectory ($\mathbf{Tr}_{t}^{0}$), by replacing the MLP for boxes with an MLP for camera poses [11]. All the downsampling ratios are aligned with the adopted 3D VAE [43], i.e., with $8n$ or $18n+$ 1 as inputs, and $2n$ or $12n+$ 1 as outputs, respectively.
4.3 Progressive Bootstrap Training
To expedite model convergence during training, we schedule the training data based on the duration of single training iterations. Specifically, we adopt a three-stage training approach: initially using low-resolution images for bootstrap training, transitioning to low-resolution short videos, and ultimately employing high-resolution long video training.

### v2

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is vital for applications like autonomous driving. Although DiT with 3D VAE has become a standard framework for video generation, it introduces challenges in controllable driving video generation, especially for geometry control, rendering existing control methods ineffective. To address these issues, we propose MagicDriveDiT, a novel approach that integrates the MVDiT block and spatial-temporal conditional encoding to enable multi-view video generation and precise geometric control. Additionally, we introduce an efficient method for obtaining contextual descriptions for videos to support diverse textual control, along with a progressive training strategy using mixed video data to enhance training efficiency and generalizability. Consequently, MagicDriveDiT enables multi-view driving video synthesis with $\times3$ resolution and 4×4$\times$4 × frame count (compared to current SOTA), rich contextual control, and geometric controls. Extensive experiments demonstrate MagicDriveDiT’s ability, unlocking broader applications in autonomous driving. Webpage: flymin.github.io/magicdrivedit/
@@ 1 Introduction
The rapid advancements in diffusion models and controllable generation has substantially broadened its range of applications [19, 37]. Especially for autonomous driving, controllable generation of high-resolution and long videos has become a critical research focus [26, 41]. High resolution facilitates the identification of fine details [21], while long videos provide richer interaction for evaluating and improving algorithms [41]. These capabilities are indispensable for enhancing the performance and reliability of autonomous systems [11, 35]. However, due to limitations in model scalability and the compression capability of VAEs [30, 2], existing methods [12, 11] are significantly constrained by both resolution and frame count, as shown in Figure 1.
@@
The introduction of DiT [29, 46] and 3D VAE [46, 42] has substantially alleviated these challenges. 3D VAE reduces computational overhead by performing spatial-temporal compression, effectively decreasing the computational cost of video generation by an order of magnitude. DiT enables quality improvements in generation with more model parameters and extensive data. Consequently, many text-to-video methods leverage DiT with 3D VAE for high-quality generation [46, 42]. However, autonomous driving requires multi-view generation with geometry control. How to effectively leverage the DiT with 3D VAE to achieve such controllable generation remains an open problem.
One of the most critical issue from controllable generation in autonomous driving is geometry control, which involves the positioning of objects and the road structures. These geometric conditions are typically time-dependent, meaning they correspond to individual video frames in a one-to-one manner. With 2D VAEs [30, 2], this characteristic allows control methods designed for images [43] to be seamlessly extended to video generation [11]. However, 3D VAE [46, 42] disrupts this alignment, as illustrated in Figure 2. The spatial-temporal latent of 3D VAE requires that the control conditions also encapsulate corresponding spatial-temporal information. Therefore, it is necessary to redesign the geometry control module to accommodate the DiT with 3D VAE framework for driving video generation.
@@
In this paper, we present MagicDriveDiT, a novel approach based on the DiT architecture, designed to tackle the challenges of high-resolution and long street-view video synthesis with precise control. Building upon DiT [29, 46] integrated with a 3D VAE for geometric control, we introduce an MVDiT block for multi-view video generation, and a novel spatial-temporal conditional encoding for latents derived from CogVAE [42], enabling precise management of geometric controls. Additionally, we enhance contextual description supports by efficiently generating captions for each video. To train the model efficiently, we employ a progressive training strategy, transitioning from short to long videos, which helps the model capture intricate details and generalize to complex scenarios. Furthermore, we utilize videos of diverse resolutions and durations to enhance the model’s generalization capability, enabling it to synthesize longer videos than those used in training.
Our MagicDriveDiT excels in generating highly realistic videos that align with text, road maps, 3D bounding boxes, and different camera perspectives, achieving significantly higher resolution and more frames than previous works (detailed in Table 1). Experiments and comparisons demonstrate the effectiveness of our training and control methods, significantly improving controllable street-view video synthesis. MagicDriveDiT’s flexibility in handling various resolutions, frames, and control signals enables the creation of novel street views suitable for simulations, expanding its potential applications across diverse domains.
@@
  • We design an efficient framework, MagicDriveDiT, leveraging progressive training to achieve high-quality, high-resolution, and long video generation (up to $848\times 1600\times 6$, 241 frames), significantly superior to prior works.
  • For geometric controls, novel spatial-temporal control for objects, road semantics, and camera trajectories is proposed while maintaining multi-frame, multi-view consistency. For text control, we enrich the contextual description on driving datasets to support diverse generation.
@@
Figure 3: Architecture Overview of MagicDriveDiT. To incorporate different conditions for video generation, MagicDriveDiT adopts a two-branch architecture as in [43] with basic STDiT3 blocks from [6, 46]. We propose the MVDiT block for multi-view consistency and Spatial-Temporal (Box/Traj.) Encoder to inject condition into the Spatial-Temporal (ST) Latent.
@@ 2 Related Work
Video Generation in Autonomous Driving. Video generation is crucial for autonomous driving, with applications in training perception models [11], testing [41], and scene reconstruction [10, 44]. It demands extensive fields of view and dynamic object motion handling, requiring precise controllability [11] and high-resolution video production [21] with more frames [10] and multiple camera perspectives [44]. Diffusion models have improved controllable multi-view video generation, yet existing models [38, 39, 45, 26, 18] lack adequate resolution and frame count for policy testing [16] and data engine applications [11, 10, 44]. By achieving unprecedented resolution and frame count, as compared in Figure 1 and Table 1, MagicDriveDiT addresses the limitations of previous driving video generation methods, enabling more potential applications.
@@
MagicDriveDiT 848$\times$1600$\times$6 241
Table 1: Comparison of Resolution and Frame Count. We only consider a single inference, since rollout notably degrades quality (see Appendix I). Only support text & image(s) conditions.
Diffusion Models and DiT Architectures. Diffusion models [13, 31, 47] generate data by learning the denoising steps from Gaussian noises to samples, widely used in image [8, 30, 37] and video [14] generation. From the modeling’s perspective, Flow Matching [22, 9] simplifies diffusion models, enhancing the efficiency of training and inference. From the implementation’s perspective, the architecture shifts from UNet [13] to DiT, due to better scalability [29, 6], especially for high-resolution tasks [5]. Our MagicDriveDiT also leverages Flow Matching and DiT for scaling to high-resolution and long video generation.
Conditional Generation. Conditional generation is crucial for various applications utilizing generative models. Cross-attention layers from LDM [30] and ControlNet’s [43] additive encoding for grid-shaped control signals are leading methods in controllable diffusion-based generation. In the street-view generation, MagicDrive [11] and MagicDrive3D [10] integrate 3D bounding boxes, BEV maps, ego trajectories, and camera poses for multi-view street scene synthesis. However, these methods are limited to spatial encoding and are not directly applicable to spatial-temporal latents from 3D VAE [42], as shown in Figure 2. MagicDriveDiT presents a novel control paradigm for spatial-temporally compressed latents from 3D VAE, enabling the controllable generation of high-resolution long videos.
@@ 3 Preliminary
Problem Formulation. This paper addresses controllable high-resolution and long video generation for street views. Given a sequence of frame descriptions $\{\mathbf{S}_{t}\},t\in\{0,\dots,T\}$, the goal is to generate corresponding street-view videos from latent variables $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$, i.e., $\{\mathbf{I}_{c,t}\}=\mathcal{G}(\{\mathbf{S}_{t}\},\mathbf{z})$, where $c\in\{0,\dots,C\}$ denotes the $C$ camera views. Here, high resolution indicates that $\mathbf{I}$ has a high resolution, and long video implies that $T$ is large.
To describe a street-view video, we adopt the conditions outlined in [10, 11]. Specifically, frame descriptions $\mathbf{S}_{t}=\{\mathbf{C},\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L},\mathbf{Tr}_{t}^{0}\}$ include the camera poses $\{\mathbf{C}_{c}\}=[\mathbf{R}_{c},\mathbf{t}_{c}]$ [^1] Typically, cameras are fixed throughout the video., a road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ representing a $w\times h$ meter road area in BEV with $c$ semantic classes, 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ where each object is described by a box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$, text $\mathbf{L}$ adding information for the whole video (e.g., weather and time of day), and the ego vehicle trajectory $\mathbf{Tr}_{t}^{0}$ which describes the transformation from the LiDAR coordinate of each frame to the first frame, i.e., $\mathbf{Tr}^{0}_{t}=[\mathbf{R}_{t}^{0},\mathbf{t}_{t}^{0}]$. All geometric data, except $\mathbf{Tr}_{t}^{0}$, is parameterized in the ego car’s local LiDAR frame.
@@ 4.1 Overview of MagicDriveDiT
Figure 3 illustrates an overview of the model architecture. Based on the STDiT-3 blocks Zheng et al. [46], we introduce Multi-View DiT (MVDiT) block for multi-view generation and employ cross-attention [30] for text, boxes, camera views, and trajectories, alongside an addictive branch [43] for maps to inject control signals.
@@
Besides, MagicDriveDiT enriches text control beyond previous works (e.g[11]) by generating contextual description through image captioning (Section 4.4). Furthermore, MagicDriveDiT employs progressive training from images to high-resolution short videos, then to long videos, enhancing diffusion model convergence. Using variable-length, multi-resolution videos in training also enables image and video generation at diverse resolutions and frame extrapolation beyond training settings (Section 4.5).
4.2 Architecture for High-Resolution Long Video
Training high-resolution, long-video diffusion models is computationally intensive, requiring significant GPU memory. DiT and 3D VAE are key to scaling these models. As noted by Peebles and Xie [29], DiT’s lower Gflops correlate with better FID, making it more efficient than UNet [11]. For memory, 3D VAE [42] compresses temporal information by 4×4$\times$4 ×, achieving a 256×256$\times$256 × compression ratio [^2] We ignore latent dimension discrepancies., reducing sequence length and memory use, especially for transformers [7]. Architectural unification [34] enables DiT to leverage advanced parallelization, e.g., sequence parallelism [46], overcoming single GPU memory limits for higher resolution and longer videos.
As depicted in Figure 3, MagicDriveDiT utilizes STDiT-3 blocks as per Zheng et al. [46]. The architecture design incorporates two significant modifications. First, to facilitate multi-view generation, the Multi-View DiT (MVDiT) block integrates a cross-view attention layer [11], as shown on the left of Figure 3. Second, given the need for handling multiple control elements, MagicDriveDiT employs cross-attention [30] for text, boxes, camera views, and trajectories, alongside an addictive branch [43] for maps to inject control signals.
4.3 Spatial-Temporal Control on 3D VAE
A major challenge with 3D VAE adoption is geometric control. As in Figure 2, geometric control manages per-frame content spatially. Previously with 2D VAE, the temporal axis persists in geometric descriptors $\{\mathbf{S}_{t}\},t\in\{1,...,T\}$, video latent and the decoded video. This temporally aligning feature enables video geometric control with spatial control methods similar to that on images [27]. However, 3D VAE yields $T/f$ latents (where $f$ is the temporal compression ratio), misaligning control signals with latents and rendering previous control techniques [11, 39] ineffective.
Besides, the issue of temporal misalignment does not exist in ambiguous control conditions (e.g., text), as these conditions are typically employed to describe the entire video. Specifically, the temporal dimension of such conditions is $1$, and they are repeated for $T/f$ times to control the entire video. Consequently, a naive approach to handle such dimension misalignment would be to follow the method of text control, where the temporal dimension is first reduced to $1$ through a temporal encoder and then repeated to $T/f$ for control. However, our experiments reveal that this operation results in trailing issues (as ”Box Reduce” in Figure 10), which we hypothesize are caused by the repeat operation. Therefore, we design a downsampling method that aligns the dimensions while preserving the distinctiveness of temporal information, thereby enabling precise control of geometric information for each frame.
@@ 4.3 Spatial-Temporal Control on 3D VAE
Our spatial-temporal encoding modules are illustrated in Figure 4. This involves re-aligning the maps ($\mathbf{M}_{t}$), boxes ($\mathbf{B}_{t}$), and trajectory ($\mathbf{Tr}_{t}^{0}$) within the scene descriptor ($\mathbf{S}_{t}$). Maps, represented as grid data, are straightforward to manage. By extending ControlNet’s [43] design, we utilize temporal downsampling modules with new trainable parameters in 3D VAE to align features between control and base blocks, as depicted in Figure 4(a).
@@
For 3D boxes, padding is applied to invisible boxes to maintain consistent box sequence length across views and frames, illustrated in Figure 3 right. We employ a downsampling module with a temporal transformer and RoPE [32] to capture temporal correlation, creating spatial-temporal embeddings aligned with video latents, as in Figure 4(b). The spatial-temporal encoder for boxes can also be adapted for ego trajectory ($\mathbf{Tr}_{t}^{0}$), by replacing the MLP for boxes with an MLP for camera poses [11].
All the downsampling ratios are aligned with the adopted 3D VAE [42], i.e., with $8n$ or $18n+$ 1 as inputs, and $2n$ or $12n+$ 1 as outputs, respectively. Section 5.3 holds more ablation studies on different downsampling methods.
Figure 5: MagicDriveDiT supports more diverse text control by enriching the textual description of driving datasets (e.g., nuScene [4], Waymo [33]) than previous works (e.g., [11]). Please see the full videos on our project page.
4.4 Enrich Text Control by Image Caption
Although geometric descriptors offer precise control over objects and maps, textual descriptions in autonomous driving datasets (e.g., nuScenes) are often restricted to basic weather (sunny/rain) and time (day/night), lacking richer scene context such as road types (e.g., highway/urban) or background elements (e.g., buildings, trees). To enhance textual control, we regenerated video captions using a Multimodal Large Language Model (MLLM) [17]. By prompting MLLM to focus on aspects beyond object categories, trajectories, and road structures, we ensure efficiency and consistency. Specifically, first, it eliminates the need for the MLLM to describe dynamics, thus only the middle frame of each video is used; second, it avoids potential conflicts between multiple control signals. Figure 5 shows the diverse contextual description supports of MagicDriveDiT.
4.5 Progressive Training with Mixed Video Types